# MagicDrive3D: Controllable 3D Generation for Any-View Rendering in Street Scenes

### v3

@@ 5.3 Ablation Study
Ablation on Enhanced Gaussian Splatting. As detailed in Section 4.3, three enhancements - prior, modeling, and loss - have been made to 3DGS. To evaluate their efficacy, each was ablated from the final algorithm, the results of which are shown in Table 4. Notations “w/o depth scale opt.” and “w/o depth opt.” represent the absence of GS loss optimization for $(s_{c,t},b_{c,t})$ and use of direct output from the monocular depth model, respectively. Each component’s removal lowered the method’s performance, while incorrect depth sometimes performs worse than the 3DGS baseline. Removal of AE in “vary-t” led to inferior PSNR but improved LPIPS, which is reasonable because AE mitigates the pixel-wise color constraint during reconstruction.
5.4 Application
Training Support for Perception Tasks. We demonstrate an application wherein street scene generation serves as a data engine for perception tasks, leveraging the advantage of any-view rendering ability of our method to improve viewpoint robustness for perception models [14]. We employ CVT [39] and the BEV segmentation task following the evaluation protocols of [39, 9]. By incorporating 4 different rigs on the FRONT camera and adding rendered views for training, the negative impact from viewpoint changes is alleviated (Table 5), exemplifying the utility of street scene generation in training perception tasks.
Render Object-level Dynamic. MagicDrive3D is capable of generating 3DGS representations of scenes, which facilitates various applications in scene editing. Our method utilizes metric scale modeling, guaranteeing that any modifications made to the scenes accurately reflect real-world physical distances. As illustrated in Figure 6, we segmented the generated GS and repositioned the object on the right. The resulting scene GS supports rendering effectively.
@@ 6 Conclusion and Discussion
This paper introduces MagicDrive3D, a unique 3D street scene generation framework that integrates geometry-free view synthesis and geometry-focused 3D representations. MagicDrive3D significantly reduces data requirements, enabling training on typical autonomous driving datasets, such as nuScenes. MagicDrive3D employs a video generation model to enhance inter-frame consistency, while the enhanced deformable GS improves reconstruction quality from generated views. Comprehensive experiments demonstrate that MagicDrive3D can produce high-quality 3D street scenes with multi-level controls. Additionally, we show that scene generation can serve as a data engine for perception tasks such as BEV segmentation.
Supplementary Material (Appendix)
@@
A More Implementation Details
For video generation, we train our generator based on the pre-trained street view image generation model from [9]. By adding the proposed relative pose control, we train 4 epochs (77040 steps) on the nuScenes training set with a learning rate of $8e^{-5}$. We follow the settings for 7-frame videos described in [9], using $224\times 400$ for each view but extending to $T=16$ frames. Consequently, for reconstruction, we select $t=8$ as the canonical space. Except we change the first 500 steps to optimize $(s_{c,t},b_{c,t})$ for each view and $\lambda_{\text{reg}_{o}}=1.0$, other settings are the same as 3DGS.
@@ A More Implementation Details
Since GS only supports perspective rendering, to stitch the view for panorama, we use code from perspective to equirectangular transformation provided by https://github.com/timy90022/Perspective-and-Equirectangular.
All our experiments are conducted with NVIDIA V100 32GB GPUs. The generation of a single scene takes about 2 minutes for video generation and about 30 minutes for deformable GS reconstruction. For reference, 3DGS reconstruction typically takes about 23 minutes for scenes of similar scales. Therefore, the proposed enhancement is efficient. As for rendering, there is no additional computation for our method compared with 3DGS.
B More Ablation Study
Ablation on Offset Choice for Deformable GS. In addition to the overall module ablation, we observe that for Deformable GS, beyond the Gaussians’ center coordinates, their attributes (including anisotropic covariance, opacity, and harmonic coefficients) can also be utilized to address local inconsistencies. To verify the effects of different choices, we randomly select 10 scenes from the nuScenes validation set for experimentation. As shown in Table I, the results obtained by adding offsets to the center coordinates (xyz) are the best. This aligns with our observation that local inconsistencies in the generated views occur primarily in the shape of objects, and thus, xyz displacements can most effectively resolve these inconsistencies.
@@ B More Ablation Study
Table I: Ablating comparison with offsets on anisotropic covariance (Cov.), opacity, and harmonic coefficients (Features) properties in GS. We randomly sample 10 scenes from the nuScenes validation set for experiments and apply color correction (cc) to all the renderings.
C More Reconstruction Baseline
Focusing on dynamic scenes, 4DGS [33] introduces comprehensive improvements over 3DGS and achieves notable results. Therefore, we replace deformable GS with 4DGS. As shown in Table II, by incorporating non-rigid dynamics, 4DGS already performs better than 3DGS. However, in our task, 4DGS underperforms compared to our deformable GS. Based on the results in Table I, we hypothesize that our reconstruction algorithm only needs to address local dynamics caused by content inconsistency. Allocating excessive dynamics at the representational level may hinder model convergence, thus 4DGS does not yield better results in our scenarios.
@@ C More Reconstruction Baseline
Table II: Comparison with 4DGS [33]. We randomly sample 10 scenes from the nuScenes validation set for experiments and apply color correction (cc) to all the renderings.
D Limitation and Future Work
As a data-centric method, MagicDrive3D sometimes struggles to generate complex objects like pedestrians, whose appearances are intricate. Additionally, areas with much texture detail (e.g., road fences) or small spatial features (e.g., light poles) are occasionally poorly generated due to limitations in the reconstruction method. Future work may focus on addressing these challenges and further improving the quality and robustness of generated 3D scenes.
E Comparison with Simple Baselines
@@
Figure I: Comparison with two baselines (LucidDreamer [6] and WonderJourney [36]) and direct stitching real images.
As shown in Figure I, we further compare MagicDrive3D with two baselines, i.e., LucidDreamer [6] and WonderJourney [36]. The former method has been proposed recently and takes text description as the only condition. Thus, it is hard to generate photo-realistic street scenes. When providing multi-view video frames from nuScenes with known camera poses, their pipeline fails to reconstruct. We suppose the reason is limited overlaps and errors from depth estimation. As suggested by the released code, we changed the image generation model to lllyasviel/control_v11p_sd15_inpaint for inpainting by providing a nuScenes image, i.e., Figure 1(a). However, due to the lack of controllability, the results from LucidDream (e.g., Figure 1(b)) are unsatisfactory. On the other hand, due to the lack of control over objects within the scene, WonderJourney struggles to generate coherent scenes. Inpainting-based methods like the two above exhibit a pronounced sense of patches and face significant challenges in achieving 360 coverage.
Figure 1(d) further shows directly stitching real data. It is also bad due to the limited overlaps between views. On the contrary, the scene generated from MagicDrive3D can render continuous panorama, as shown in Figure 1(e), which is also controllable through multiple conditions.
Note that, panorama generation is only one of the applications of our generated scenes. We show them just for convenient qualitative comparison within the paper. Since our scene generation contains geometric information, they can be rendered from any camera view, as shown in Figure 5.
F Implementation Detail of Appearance Embedding
We show in Figure II the detailed architecture of the CNN used in our appearance modeling. The AE map is 32×32$\times$32 × smaller than the input image to reduce the computational cost. Hence, we first downsample the input image by 32×32$\times$32 ×. Then, we use $3\times 3$ convolution for feature extraction and pixel shuffle for upsampling. Each convolution layer is activated by ReLU.
Figure II: The CNN architecture of appearance modeling, as introduced in Section 4.3.
G Broader Impacts
@@
H More Qualitative Results
We show more generated street scenes from MagicDrive3D in Figure III and Figure IV.
Figure III: Generated street scenes from MagicDrive3D. We adopt control signals from the nuScenes validation set. We crop the center part for better visualization.
Figure IV: Generated street scenes from MagicDrive3D. We adopt control signals from nuScenes validation set. The black regions are not fully covered, constrained by the camera’s FOV.

### v4

@@ 5.5 Ablation Study
Effectiveness of the Enhanced GS pipeline. We first show the effectiveness of the proposed enhanced GS pipeline through an experimental setting similar to the evaluation of a general reconstruction method [13], i.e., comparing renderings with ground truth (GT) images, with 3DGS [13] as the baseline method. Specifically, two testing scenarios are employed (as shown in Table I, appendix) : 360, where all six views from $t=9$ are reserved for testing; and vary-t, where one view is randomly sampled from different $t$ to assess long-range reconstruction ability through $t$. Here, the views generated by the video generator are treated as GT. Performance is assessed using L1, PSNR, SSIM [34], and LPIPS [43], as per Kerbl et al. [13].
Among all the metrics in Table 4, our enhanced GS not only improves reconstruction quality for training views but also drastically enhances generalization quality for testing views, compared to 3DGS. We also include more comparison with 4DGS [37] as another baseline in Appendix E.
Ablation on Enhanced Gaussian Splatting. As detailed in Section 4.3, three enhancements—prior, modeling, and loss—were made to 3DGS. To evaluate their efficacy, each was ablated from the final algorithm, with results shown in Table 5. Notations “w/o depth scale opt.” and “w/o depth opt.” indicate the absence of GS loss optimization for $(s_{c,t},b_{c,t})$ and the use of direct output from the monocular depth model, respectively. Removing each component reduced performance, and incorrect depth sometimes performed worse than the 3DGS baseline. Removing AE in “vary-t” led to lower PSNR but improved LPIPS, as AE mitigates pixel-wise color constraints during 3DGS generation.
Appendix D contains more ablation studies.
@@ 6 Conclusion and Discussion
This paper introduces MagicDrive3D, a novel framework for controllable 3D street scene generation that integrates multi-view video synthesis with 3D representations. MagicDrive3D uses a relative pose embedding in video generation to enhance inter-frame quality while providing camera pose data for each view. The enhanced fault-tolerant GS in MagicDrive3D improves 3DGS generation quality from the generated camera views. Consequently, MagicDrive3D significantly reduces data collection requirements, enabling training on standard autonomous driving datasets such as nuScenes. Comprehensive experiments demonstrate that MagicDrive3D can produce high-quality 3D street scenes with multi-level controls. Furthermore, the generated scenes enhance viewpoint robustness in perception tasks like BEV segmentation.
Appendix (Supplementary Material)
@@
Appendix A Further Discussion on Differences from Reconstruction Works
The main focus of this work is the generation of non-existent 3D scenes, which distinguishes it from the reconstruction of existing scenes (such as [39, 48]). While 3D scene reconstruction also leverages video generation models to enhance quality [22, 40], its goal is to produce novel views of the same scene from a given set of camera angles. In contrast, our model can directly generate entirely new 3D scenes through latent variable sampling after training. The key difference is that our approach enables the creation of entirely new scenes, a capability that reconstruction methods cannot achieve (as shown in Table 1). Specifically, our model can generate novel scenes that, while adhering to control conditions, allow for the updating of objects and environmental features, such as buildings, across different samples. Figures 1 and 6 demonstrate our method’s ability to construct diverse 3D scenes based on varying conditions. We hope this discussion helps readers gain a clearer understanding of the contribution of this work.
Appendix B More Implementation Details
For video generation, we train our generator based on the pre-trained street view image generation model from Gao et al. [11]. By adding the proposed relative pose control, we train 4 epochs (77040 steps) on the nuScenes training set with a learning rate of $8e^{-5}$. We follow the settings for 7-frame videos described in Gao et al. [11], using 224×400224$\times$ 400224 × 400 for each view but extending to $T=16$ frames. Consequently, for 3DGS generation, we select $t=8$ as the canonical space. Except we change the first 500 steps to optimize $(s_{c,t},b_{c,t})$ for each view and $\lambda_{\text{reg}_{o}}=1.0$, other settings are the same as 3DGS.
@@ Appendix B More Implementation Details
Since GS only supports perspective rendering, to stitch the view for panorama, we use code provided by https://github.com/timy90022/Perspective-and-Equirectangular to transform from perspective to equirectangular.
All our experiments are conducted with NVIDIA V100 32GB GPUs. The generation of a single scene takes about 2 minutes for video generation and 30 minutes for GS generation. For reference, 3DGS reconstruction typically takes about 23 minutes for scenes of similar scales. Therefore, the proposed enhancement is efficient. As for rendering, there is no additional computation for our method compared with 3DGS.
As mentioned in Section 5.5, we use two testing scenarios to show the effectiveness of the enhanced GS pipeline as an ablation study, i.e., 360 and vary-t. We use one testing trajectory as an example to illustrate the training/testing view splits in Table I.
  name #test #train camera poses
  360 6 90 [Uncaptioned image]
vary-t 12 84 [Uncaptioned image]
Table I: Two settings for ablation study on the enhanced GS pipeline. Testing views are in green while training views are in red.
Appendix C Optimization Flow
We demonstrate the overall optimization flow of the proposed FTGS in Algorithm 1. Line 2 is the first optimization of monocular depths. Lines 4-8 refer to the second optimization of the monocular depths. Lines 10-16 are the main loop for FTGS, where we consider temporal offsets on Gaussians, camera pose optimization for local inconsistency, and AEs for appearance discrepancies among views.
Algorithm 1 Enhanced Fault-Tolerant Gaussian Splatting (FTGS)
0: camera views $\{\mathcal{I}_{i}\}$, camera parameters $\{\mathbf{P}^{0}_{i}\}$, monocular depth $\{\mathcal{D}_{i}\}$, optimization steps for depth $s_{D}$, camera pose $s_{\text{C}}$, and GS $s_{\text{GS}}$
0: FTGS of the scene $\{\boldsymbol{\mu}_{p},\boldsymbol{\mu}^{o}_{p},\boldsymbol{\Sigma}_{p},\text{SH}_{p}\}$, and optimized camera pose $\{\mathbf{P}^{0}_{i}\}$
[…]
2: Optimize $(s_{c,t},b_{c,t})$ with $\mathcal{P}_{\text{SfM}}$ for each $\{c,t\}$
3: Random initialize AEs $\{\mathbf{e}_{i}\}$
4:for step in $1,\dots,s_{D}$ do
5:  Random pick one view $\mathcal{I}_{i}$
6:  $\mathcal{L}=\mathcal{L}_{\text{AEGS}}(\mathcal{I}_{i},\mathcal{I}^{r}_{i},\mathbf{e}_{i})$
7:  Update $(s,b),\mathbf{e}_{i},\boldsymbol{\Sigma},\text{SH}$ with $\nabla\mathcal{L}$
8:end for
[…]
10:for step in $s_{D},\dots,s_{\text{GS}}$ do
11:  Random pick one view $\mathcal{I}_{i}$ and get its $t$
12:  $\mathcal{L}=\mathcal{L}_{\text{FTGS}}(\mathcal{I}_{i},\mathcal{I}^{r}_{i},\mathbf{e}_{i},\boldsymbol{\mu}^{o}(t))$
13:  Update $\boldsymbol{\mu},\boldsymbol{\mu}^{o}(t),\mathbf{e}_{i},\boldsymbol{\Sigma},\text{SH}$ with $\nabla\mathcal{L}$
14:  if step $>s_{\text{C}}$ then
15:   Update $\mathbf{P}^{0}_{i}$ with $\nabla\mathcal{L}$
16:  end if
17:end for
Table II: Ablating comparison with offsets on anisotropic covariance (Cov.), opacity, and harmonic coefficients (Features) properties in GS. We randomly sample 10 scenes from the nuScenes validation set for experiments and apply color correction (cc) to all the renderings.
@@ Appendix C Optimization Flow
Appendix D More Ablation Study
Ablation on Offset Choice for Fault-Tolerant GS. In addition to the overall module ablation, we observe that for Fault-Tolerant GS, beyond the Gaussians’ center coordinates, their attributes (including anisotropic covariance, opacity, and harmonic coefficients) can also be utilized to address local inconsistencies. To verify the effects of different choices, we randomly select 10 scenes from the nuScenes validation set for experimentation. As shown in Table II, the results obtained by adding offsets to the center coordinates (xyz) are the best. This aligns with our observation that local inconsistencies in the generated views occur primarily in the shape of objects, and thus, xyz displacements can most effectively resolve these inconsistencies.
Qualitative Comparison with 3DGS baseline. Figure I shows the results of reconstructing the generated video directly using the 3DGS algorithm. As discussed in Section 4.3, 3DGS tends to amplify artifacts in the generated view, whereas our improved FTGS pipeline effectively controls the impact of these artifacts, ensuring the quality of the 3D scene generation.
Figure I: Qualitative Comparison with 3DGS baseline on the same scene shown in Figure 5. With our improved FTGS pipeline, the generated 3DGS contains fewer floaters, leading to higher-quality rendering.
Appendix E More Reconstruction Baseline
Focusing on dynamic scenes, 4DGS [37] introduces comprehensive improvements over 3DGS and achieves notable results. Therefore, we replace Fault-Tolerant GS with 4DGS. As shown in Table III, by incorporating non-rigid dynamics, 4DGS already performs better than 3DGS. However, in our task, 4DGS underperforms compared to our Fault-Tolerant GS. Based on the results in Table II, we hypothesize that our FTGS algorithm only needs to address local inconsistency caused by content inconsistency. Allocating excessive degrees-of-freedom at the representational level may hinder model convergence, thus 4DGS does not yield better results in our scenarios.
Table III: Comparison with 4DGS [37]. We randomly sample 10 scenes from the nuScenes validation set for experiments and apply color correction (cc) to all the renderings.
@@ Appendix E More Reconstruction Baseline
Appendix F Limitation and Future Work
As a data-centric method, MagicDrive3D sometimes struggles to generate complex objects like pedestrians, whose appearances are intricate. Additionally, areas with much texture detail (e.g., road fences) or small spatial features (e.g., light poles) are occasionally poorly generated due to limitations in the 3DGS method. Future work may focus on addressing these challenges and further improving the quality and robustness of generated 3D scenes.
Appendix G Comparison with Simple Baselines
@@
Figure II: Comparison with two baselines (LucidDreamer [7] and WonderJourney [42]) and direct stitching real images.
As shown in Figure II, we further compare MagicDrive3D with two baselines, i.e., LucidDreamer [7] and WonderJourney [42]. The former method has been proposed recently and takes text description as the only condition. Thus, it is hard to generate photo-realistic street scenes. When providing multi-view video frames from nuScenes with known camera poses, their pipeline fails to reconstruct. We suppose the reason is limited overlaps and errors from depth estimation. As suggested by the released code, we changed the image generation model to lllyasviel/control_v11p_sd15_inpaint for inpainting by providing a nuScenes image, i.e., Figure 2(a). However, due to the lack of controllability, the results from LucidDream (e.g., Figure 2(b)) are unsatisfactory. On the other hand, due to the lack of control over objects within the scene, WonderJourney struggles to generate coherent scenes. Inpainting-based methods like the two above exhibit a pronounced sense of patches and face significant challenges in achieving 360 coverage.
Figure 2(d) further shows directly stitching real data. It is also bad due to the limited overlaps between views. On the contrary, the scene generated from MagicDrive3D can render a continuous panorama, as shown in Figure 2(e), which is also controllable through multiple conditions.
Note that, panorama generation is only one of the applications of our generated scenes. We show them just for convenient qualitative comparison within the paper. Since our scene generation contains geometric information, it can be rendered from any camera view, as shown in Figure 5.
Appendix H Implementation Detail of Appearance Embedding
We show in Figure III the detailed architecture of the CNN used in our appearance modeling. The AE map is 32×32$\times$32 × smaller than the input image to reduce the computational cost. Hence, we first downsample the input image by 32×32$\times$32 ×. Then, we use $\times 3$ 3 convolution for feature extraction and pixel shuffle for upsampling. Each convolution layer is activated by ReLU.
Figure III: The CNN architecture of appearance modeling, as introduced in Section 4.3.
Appendix I Broader Impacts
@@
Appendix J More Qualitative Results
We show more generated street scenes from MagicDrive3D in Figure IV and Figure V.
Figure IV: Generated street scenes from MagicDrive3D. We adopt control signals from the nuScenes validation set. We crop the center part for better visualization.
Figure V: Generated street scenes from MagicDrive3D. We adopt control signals from nuScenes validation set. The black regions are not fully covered, constrained by the camera’s FOV.